# DoGCLR: Dominance-Game Contrastive Learning Network for Skeleton-Based Action Recognition

### v1

@@ III-A Overview of the DoGCLR Network
Let $\Gamma=(N,(C_{i})_{i\in N},(u_{i})_{i\in N})$ denote a game, where $N$ represents the set of players, $C_{i}$ represents the set of strategies of player $i$, and $u_{i}$ represents the payoff function of player $i$. If there exists a strategy profile $c=(c_{1},\cdots,c_{|N|})$, where $c_{i}\in C_{i}$, and $|N|$ denotes the number of elements in the set $N$, such that for any player $i\in N$, any $d_{i}\in C_{i}$, $u_{i}(c)>u_{i}(d)$, where $d=(d_{1},\cdots,d_{|N|})$, then $c$ is called a dominance strategy profile, and the game is referred to as a Dominance Game (DG).
@@
Specifically, in the DW-KRM, the Global Statistical Benchmark Pose (GSBP) is computed first. The cosine similarity between the GSBP and the input skeleton data is then calculated, and the negative gradient of the node features with respect to this similarity yields the Discrepancy-Degree (DD) of each node. Subsequently, in the Joint-Degree Activation Module (JDAM), the adjacency matrix of the skeleton graph is used to count the number of edges connected to each node, which is then normalized to obtain the Joint-Degree (JD). Combining DD and JD produces the spatio-temporal composite weight, which, together with the node key features, locates the key motion regions that determine the fundamental properties of actions. In the DGA module, based on the key motion regions located by DW-KRM, style transfer is first applied to nodes with lower JD values. Then, from both similarity and dissimilarity perspectives, different levels of data augmentation are applied to the key and non-key motion regions respectively, enriching the diversity of samples while maintaining semantic integrity. Finally, in the EDGRQ module, an entropy-based DG is introduced, where entropy serves as the utility function and dominance strategy acts as the convergence mechanism to enhance the value of negative samples. The overall framework of DoGCLR is illustrated in Fig. 1, in which the positive sample optimization is achieved by DW-KRM and DGA, and the negative sample screening is performed by EDGRQ. Together, these components constitute the dual-dimensional dominance system of “positive sample optimization and negative sample selection”.
@@ III-B Spatio-temporal Dual-Weight synergistic localization algorithm-based Key Region partitioning Module
Figure 2: Block diagram of DW-KRM. The data and Global Statistical Benchmark Pose (GSBP) are encoded by the key encoder $f_{k}(\cdot)$ and MLP $g_{k}(\cdot)$ to compute their respective features. The Discrepancy-Degree (DD) is then fed into the Joint-Degree Activation Module (JDAM), where it is multiplied by the Joint-Degree (JD) calculated within JDAM to produce the spatio-temporal composite weight $\alpha_{c}^{(i)}$. This output from JDAM provides guidance for subsequent partitioned data augmentation. $\bigotimes$ denotes matrices multiplication.
In self-supervised contrastive learning, positive samples are required to possess high information content and strong representativeness. High information content requires positive samples to contain rich fine-grained motion information, while strong representativeness requires positive samples to accurately reflect the core features of action categories based on the key joints of the skeleton. Therefore, this paper proposes a Spatio-temporal Dual-Weight synergistic localization algorithm. After the sample is input into DW-KRM, based on the idea of local dominance, two types of weights — the Discrepancy-Degree (DD) and the Joint-Degree (JD) calculated in the Joint-Degree Activation Module (JDAM) — are fused to generate a key motion region mask, thereby achieving adaptive focus on key motion regions. The block diagram of DW-KRM is shown in Fig. 2.
To facilitate the explanation of the proposed Spatio-temporal Dual-Weight synergistic localization algorithm, the definitions of Global Statistical Benchmark Pose (GSBP), Discrepancy-Degree (DD), and Joint-Degree (JD) are introduced as follows.
@@
Let there be a skeleton dataset $\{X^{(1)},X^{(2)},\cdots,X^{(N)}\}$, where the total number of samples is $N$, and $X^{(i)}\in\mathbb{R}^{|C_{X}|\times|T|\times|V|}$, where $|C_{X}|$, $|T|$, and $|V|$ respectively denote the channel, frame, and joint set sizes of $X^{(i)}$. Then the Global Statistical Benchmark Pose (GSBP) is the algebraic mean of the dataset along the sample dimension, that is:
@@
Definition 3 (Discrepancy-Degree, DD).
@@
where $\|\cdot\|$ denotes the $\ell_{2}$-norm of the feature vector. The Discrepancy-Degree (DD) of the $(t,v)$-th node of the $i$-th sample is defined as the negative derivative of the similarity $sim(z^{(i)},\overline{z})$ with respect to the key feature $h^{(i)}_{ctv}$ on the channel of the encoded sample $X^{(i)}$:
@@
The three definitions above form the foundation of the Spatio-temporal Dual-Weight synergistic localization algorithm and provide both temporal and spatial constraints for subsequent region localization. First, the sample $X^{(i)}$ and the GSBP $\overline{X}$ are input into the key encoder and the MLP to obtain feature $z^{(i)}$ and $\overline{z}$. The DD is then computed by Eq. (3). Next, the calculated DD is input into JDAM, and the degree-based weight vector $w_{degree}$ is computed according to Eq. (5). By combining DD and JD, the spatio-temporal composite weight $\alpha_{c}^{(i)}$ is defined to measure the importance of neuron $c$, as shown in:
@@
From Eq. (6), the spatio-temporal composite weight $\alpha_{c}^{(i)}$ considers both DD ($\Delta h_{ctv}^{(i)}$) and JD ($w_{degree}[v]$), exploring nodes with higher DD from the temporal perspective and focusing on nodes with higher JD in the spatial topology of the skeleton graph. Thus, the distribution of weights across key regions is determined collaboratively by the two factors DD and JD.
From the perspective of game theory, DW-KRM can be regarded as a local DG. In this game, each skeleton node acts as a player whose strategy is how to allocate attention weights between DD ($\Delta h_{ctv}^{(i)}$) in the temporal dimension and JD ($w_{degree}[v]$) in the spatial dimension to maximize its information gain in the global representation. The spatio-temporal composite weight $\alpha_{c}^{(i)}$ of a node can be regarded as its payoff function $\alpha_{c}^{(i)}=f(\Delta h_{ctv}^{(i)},w_{degree}[v])$, and the competition and cooperation among all nodes over weight allocation form a non-cooperative DG system. Finally, the distribution of node weights converges to a local dominant equilibrium, meaning that no node can achieve higher payoff by adjusting its own strategy while others remain unchanged. This game-theoretic feature selection mechanism enables DW-KRM to possess adaptability and robustness in key-region extraction, thereby supporting the subsequent DGA.
@@ III-C Dual-scale-Game-based partitioned data Augmentation module
where $X_{Final}^{(i)}\in\mathbb{R}^{|T|\times|V|}$ denotes the final augmented data. $M^{(i)}=(M_{tv}^{(i)})_{|T|\times|V|}\in\mathbb{R}^{|T|\times|V|}$ represents a binary-encoded matrix. For skeleton $i$, if the key motion region mask $A_{tv}^{(i)}$ of joint $v$ in frame $t$ exceeds the importance threshold $\theta$, then the element $M_{tv}^{(i)}$ is set to $1$; otherwise $M_{tv}^{(i)}$ is set to 0. The operator $\odot$ denotes the Hadamard product, i.e., element-wise multiplication of matrices.
@@ IV-C1 Linear Evaluation Protocol
A further comparison is conducted on the PKU-MMD Part I and Part II datasets, as summarized in Table IV. The PKU-MMD dataset includes more realistic and fine-grained motion interactions, providing a complementary evaluation to the NTU benchmarks. Part I contains fewer samples and simpler motion patterns. Since the difference between the Global Statistic Benchmark Pose (GSBP) and the true GSBP depends on dataset size, a smaller dataset leads to reduced differentiation in Discrepancy-Degree (DD) among joints. Consequently, DoGCLR may show slight bias in localizing key regions, resulting in performance in Table IV that is not the highest. However, its accuracy is only 0.4% lower than the best algorithm and 7.0% higher than the baseline AimCLR, while still outperforming most existing methods and maintaining stable recognition capability on small-scale data. This suggests that the proposed dominance-guided mechanism preserves its discriminative potential even when structural cues are limited.

### v2

@@ III-A Overview of the DoGCLR Network
Let $\Gamma=(N,(C_{i})_{i\in N},(u_{i})_{i\in N})$ denote a game, where NN represents the set of players, $C_{i}$ represents the set of strategies of player $i$, and $u_{i}$ represents the payoff function of player ii. If there exists a strategy profile $c=(c_{1},\cdots,c_{|N|})$, where $c_{i}\in C_{i}$, and $|N|$ denotes the number of elements in the set NN, such that for any player $i\in N$, any $d_{i}\in C_{i}$, $u_{i}(c)>u_{i}(d)$, where $d=(d_{1},\cdots,d_{|N|})$, then $c$ is called a dominance strategy profile, and the game is referred to as a Dominance Game (DG).
@@
Specifically, in the DW-KRM, the Global Statistical Benchmark Pose (GSBP) is computed first. The cosine similarity between the GSBP and the input skeleton data is then calculated, and the negative gradient of the node features with respect to this similarity yields the Discrepancy-Degree ($D$) of each node. Subsequently, in the Joint-Degree Activation Module (JDAM), the adjacency matrix of the skeleton graph is used to count the number of edges connected to each node, which is then normalized to obtain the Joint-Degree (JD). Combining $D$ and JD produces the spatio-temporal composite weight, which, together with the node key features, locates the key motion regions that determine the fundamental properties of actions. In the DGA module, based on the key motion regions located by DW-KRM, style transfer is first applied to nodes with lower JD values. Then, from both similarity and dissimilarity perspectives, different levels of data augmentation are applied to the key and non-key motion regions respectively, enriching the diversity of samples while maintaining semantic integrity. Finally, in the EDGRQ module, an entropy-based DG is introduced, where entropy serves as the utility function and dominance strategy acts as the convergence mechanism to enhance the value of negative samples. The overall framework of DoGCLR is illustrated in Fig. 1, in which the positive sample optimization is achieved by DW-KRM and DGA, and the negative sample screening is performed by EDGRQ. Together, these components constitute the dual-dimensional dominance system of “positive sample optimization and negative sample selection”.
@@ III-B Spatio-temporal Dual-Weight synergistic localization algorithm-based Key Region partitioning Module
Figure 2: Block diagram of DW-KRM. The data and Global Statistical Benchmark Pose (GSBP) are encoded by the key encoder $f_{k}(\cdot)$ and MLP $g_{k}(\cdot)$ to compute their respective features. The Discrepancy-Degree ($D$) is then fed into the Joint-Degree Activation Module (JDAM), where it is multiplied by the Joint-Degree (JD) calculated within JDAM to produce the spatio-temporal composite weight $\alpha_{c}^{(i)}$. This output from JDAM provides guidance for subsequent partitioned data augmentation. $\bigotimes$ denotes matrices multiplication.
In self-supervised contrastive learning, positive samples are required to possess high information content and strong representativeness. High information content requires positive samples to contain rich fine-grained motion information, while strong representativeness requires positive samples to accurately reflect the core features of action categories based on the key joints of the skeleton. Therefore, this paper proposes a Spatio-temporal Dual-Weight synergistic localization algorithm. After the sample is input into DW-KRM, based on the idea of local dominance, two types of weights — the Discrepancy-Degree ($D$) and the Joint-Degree (JD) calculated in the Joint-Degree Activation Module (JDAM) — are fused to generate a key motion region mask, thereby achieving adaptive focus on key motion regions. The block diagram of DW-KRM is shown in Fig. 2.
To facilitate the explanation of the proposed Spatio-temporal Dual-Weight synergistic localization algorithm, the definitions of Global Statistical Benchmark Pose (GSBP), Discrepancy-Degree ($D$), and Joint-Degree (JD) are introduced as follows.
@@
Let there be a skeleton dataset $\{X^{(1)},X^{(2)},\cdots,X^{(N)}\}$, where the total number of samples is NN, and $X^{(i)}\in\mathbb{R}^{|C_{X}|\times|T|\times|V|}$, where $|C_{X}|$, $|T|$, and $|V|$ respectively denote the channel, frame, and joint set sizes of $X^{(i)}$. Then the Global Statistical Benchmark Pose (GSBP) is the algebraic mean of the dataset along the sample dimension, that is:
@@
Definition 3 (Discrepancy-Degree, $D$).
@@
where $\|\cdot\|$ denotes the $\ell_{2}$-norm of the feature vector. The Discrepancy-Degree ($D$) of the $(t,v)$-th node of the $i$-th sample is defined as the negative derivative of the similarity $sim(z^{(i)},\overline{z})$ with respect to the key feature $h^{(i)}_{ctv}$ on the channel of the encoded sample $X^{(i)}$:
@@
The three definitions above form the foundation of the Spatio-temporal Dual-Weight synergistic localization algorithm and provide both temporal and spatial constraints for subsequent region localization. First, the sample $X^{(i)}$ and the GSBP $\overline{X}$ are input into the key encoder and the MLP to obtain feature $z^{(i)}$ and $\overline{z}$. The $D$ is then computed by Eq. (3). Next, the calculated $D$ is input into JDAM, and the degree-based weight vector $w_{degree}$ is computed according to Eq. (5). By combining $D$ and JD, the spatio-temporal composite weight $\alpha_{c}^{(i)}$ is defined to measure the importance of neuron $c$, as shown in:
@@
From Eq. (6), the spatio-temporal composite weight $\alpha_{c}^{(i)}$ considers both $D$ ($\Delta h_{ctv}^{(i)}$) and JD ($w_{degree}[v]$), exploring nodes with higher $D$ from the temporal perspective and focusing on nodes with higher JD in the spatial topology of the skeleton graph. Thus, the distribution of weights across key regions is determined collaboratively by the two factors $D$ and JD.
From the perspective of game theory, DW-KRM can be regarded as a local DG. In this game, each skeleton node acts as a player whose strategy is how to allocate attention weights between $D$ ($\Delta h_{ctv}^{(i)}$) in the temporal dimension and JD ($w_{degree}[v]$) in the spatial dimension to maximize its information gain in the global representation. The spatio-temporal composite weight $\alpha_{c}^{(i)}$ of a node can be regarded as its payoff function $\alpha_{c}^{(i)}=f(\Delta h_{ctv}^{(i)},w_{degree}[v])$, and the competition and cooperation among all nodes over weight allocation form a non-cooperative DG system. Finally, the distribution of node weights converges to a local dominant equilibrium, meaning that no node can achieve higher payoff by adjusting its own strategy while others remain unchanged. This game-theoretic feature selection mechanism enables DW-KRM to possess adaptability and robustness in key-region extraction, thereby supporting the subsequent DGA.
@@ III-C Dual-scale-Game-based partitioned data Augmentation module
where $X_{Final}^{(i)}\in\mathbb{R}^{|T|\times|V|}$ denotes the final augmented data. $M^{(i)}=(M_{tv}^{(i)})_{|T|\times|V|}\in\mathbb{R}^{|T|\times|V|}$ represents a binary-encoded matrix. For skeleton $i$, if the key motion region mask $A_{tv}^{(i)}$ of joint $v$ in frame $t$ exceeds the importance threshold $\theta$, then the element $M_{tv}^{(i)}$ is set to 11; otherwise $M_{tv}^{(i)}$ is set to 0. The operator $\odot$ denotes the Hadamard product, i.e., element-wise multiplication of matrices.
@@ IV-C1 Linear Evaluation Protocol
A further comparison is conducted on the PKU-MMD Part I and Part II datasets, as summarized in Table IV. The PKU-MMD dataset includes more realistic and fine-grained motion interactions, providing a complementary evaluation to the NTU benchmarks. Part I contains fewer samples and simpler motion patterns. Since the difference between the Global Statistic Benchmark Pose (GSBP) and the true GSBP depends on dataset size, a smaller dataset leads to reduced differentiation in Discrepancy-Degree ($D$) among joints. Consequently, DoGCLR may show slight bias in localizing key regions, resulting in performance in Table IV that is not the highest. However, its accuracy is only 0.4% lower than the best algorithm and 7.0% higher than the baseline AimCLR, while still outperforming most existing methods and maintaining stable recognition capability on small-scale data. This suggests that the proposed dominance-guided mechanism preserves its discriminative potential even when structural cues are limited.